OpenStudy (lukecrayonz):

Find the cross product of j x k (And yes, that is all that is given, so I have no idea what I'm doing)

OpenStudy (lukecrayonz):

Here is the full picture, number two. http://screensnapr.com/e/kvvXYJ.png

OpenStudy (anonymous):

ixj = k jxk = i kxi = j

OpenStudy (agreene):

j x k = jk sin (theta) i

OpenStudy (agreene):

well, these are unit vectors so what lady in red said is true.

OpenStudy (lukecrayonz):

so jxk = i?

OpenStudy (lukecrayonz):

I feel like I need more than that for my answer.

OpenStudy (agreene):

you can do the work with the definition of cross products, but you'll find the answer to be i hat (the i unit vector)

OpenStudy (lukecrayonz):

And sketching it?

OpenStudy (agreene):

assuming it's a normal plane (right hand rule) it will be a line on the "x" axis one unit long

OpenStudy (phi):

\[\left| \begin{matrix}i &j& k \\ 0& 1&0 \\0&0&1\end{matrix} \right|=i\cdot1-j\cdot 0+k\cdot 0=i\]

OpenStudy (lukecrayonz):

Is it possible to sketch it here? I don't get what you mean by the right hand rule, and I looked at that page she gave me but it still doesn't make any sense.

OpenStudy (agreene):